OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = 3x − 4, find f[g(3)]. 11 17 5 7

jhonyy9 (jhonyy9):

first calcule g(3)=3*3 -4=9-4=5 and now calcule f(g(3))=f(5)=2*5+1=10+1=11 ok ?

OpenStudy (anonymous):

thanx